Hi, I am in need of help. I need to show all the steps I utilized to get the final answer. If someone would be so kind as to help me, as I don't know where I should get started. I would greatly appreciate it. :o)
Pg. 828, # 55
Directions: The formula S=C(1+r)^t models inflation, where C= the value today, r= the annual inflation rate, and S=the inflated value t years from now. Use this formula to solve. Round your answer to the nearest dollar.

If the inflation rate is 6%, how much will a house now worth $465,000 be worth in 10 years?

Answer by mananth(16946)

The formula S=C(1+r)^t models inflation, where C= the value today, r= the annual inflation rate, and S=the inflated value t years from now. Use this formula to solve. Round your answer to the nearest dollar.

If the inflation rate is 6%, how much will a house now worth $465,000 be worth in 10 years?
..
S=C(1+r)^t
C=465,000
r=6% , = 0.06
t=10 years
...
S = 450000(1+0.06)^10
S= 450000*(1.06)^10
S=$805,881 after 10 years
...
m.ananth@hotmail.ca